OpenStudy (anonymous):

Find the center + radius of the circle: Equation: (x+2)^2 + (y+10)^2 = 4

OpenStudy (bahrom7893):

(-2;-10) i think

OpenStudy (bahrom7893):

I'm sure agent0smith will explain why

OpenStudy (agent0smith):

\[\Large (x+2)^2 + (y+10)^2 = 4\] This is the standard form of a circle, with center (h, k) and radius r. \[\Large (x-h)^2 + (y-k)^2 = r^2\] Now compare that to yours: \[\Large \left( x-(-2) \right)^2 + \left( y- (-10)\right) ^2 = 2^2 \]

OpenStudy (bahrom7893):

basically you have an equation: (x-h)^2+(y-k)^2=r^2 and (h;k) is the center and r is the radius

OpenStudy (anonymous):

thank you guys im really not that good in math

OpenStudy (anonymous):

so it's -2 and - 10 and radius is 4?

OpenStudy (agent0smith):

Center is correct. Pay more attention to the r^2 in the equation above.

OpenStudy (agent0smith):

Center (-2, -10)

OpenStudy (anonymous):

2?

OpenStudy (agent0smith):

Correct.

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (anonymous):

:)

OpenStudy (agent0smith):

Welcome :)